OpenStudy (anonymous):

if a $5000 piece of equipment depreciates at a rate of 5% per year, how much will it be worth after 5 years?

OpenStudy (anonymous):

\[Interest = Principle \cdot Rate ^ {Time}\] Since it is depreciating by 5%, our rate is 100% - 5% = 95%. So we plug in those values: \[$5000 \cdot 0.95 ^ 5 = $3868.90\]